OpenStudy (zion):

solve for x: log sub8 15+ log sub8 14- log sub8 105= log sub8 X

OpenStudy (anonymous):

logbase8(15x14)-log base8105=logbase8 (X) log8(210/105)=log8(x) log8(2)=log8x x=2

OpenStudy (zion):

Please share to me some parts of ur brain! hahah! i totally suck in math and i really admire those people who are expert in this kind of "suffering (for me, this is a suffering)". hahaha.

OpenStudy (anonymous):

oh !!!!! are u on facebook :D

OpenStudy (zion):

yeah.! :D

OpenStudy (anonymous):

Zion, you need to know the basic laws of logarithms. \[\ln a + \ln b = \ln ab\] \[\ln a - \ln b = \ln a/b\] stuff like these.

OpenStudy (zion):

oohhhh.. thanks! it helps me alot.! :D

OpenStudy (zion):

what is ln by the way?